Double Check: If they are wrong, can you help explain why they're wrong

In garden pea plants, round seeds (R) are dominant to wrinkled seeds (r). In an ideal pea plant population exhibiting Hardy–Weinberg equilibrium, 35 pea plants out of 92 have wrinkled seeds.

Calculate the frequency of homozygous round pea plants in the population. Express your answer using two significant digits.

Answer: 15, is that correct?

_________

In mice, black fur (W) is dominant over white fur (w). In an ideal mouse population exhibiting Hardy–Weinberg equilibrium, 200 mice out of 240 have white fur.

Calculate the number of mice that are heterozygous dominant for the fur colour gene in the population.

Answer: 16

________

In the late 1400s, John Cabot and Jacques Cartier made various voyages to what is now known as) Canada. Over the next 200 years, the eastern coastal areas of North America were settled by mostly French, English, and Spanish citizens.

What form of genetic drift is represented by the movement of the first European population to North America at this time?

Select one:

a. Founder effect

b. Controlled breeding

c. Natural selection

d. Bottleneck effect

I picked b, is that correct?
Answer #1

Answer:

1).

The frequency of wrinkled seeds = q^2 = 35/92 = 0.38

Frequency of recessive allele = q = SQRT of 0.38 = 0.62

As p+q = 1

The frequency of dominant (round allele) = p = 1 – 0.62 = 0.38

The frequency of homozygous round pea plant = p^2 = 0.38 * 0.38 = 0.15 or 15%

2).

The frequency of white fur = q^2 = 200/240 = 0.83

Frequency of recessive allele = q = SQRT of 0.83 = 0.91

As p+q = 1

The frequency of dominant (round allele) = p = 1 – 0.91 = 0.09

The frequency of heterozygous fur = 2pq = 2*0.09*0.91 = 0.16 0r 16%